OpenStudy (ledah):

A rectangle is twice as long as it is wide. Its width is 4 cm. What is its area?

Parth (parthkohli):

What is the number twice as 4?

OpenStudy (ledah):

8 I'm guessing..

OpenStudy (anonymous):

the area of rectangular is A=length * width here width=4cm and as mention in the question that length is twice as its width so length=4*2=8 nw put this in above formula @Ledah hope it help

Parth (parthkohli):

Hmm. So what is 4 * 8?

Parth (parthkohli):

No need to guess. That was right ;)

OpenStudy (ledah):

ok.

OpenStudy (anonymous):

A=8*4 A=32cm^2

Parth (parthkohli):

Huh? @Qibtiya It's actually 4 * 8

Parth (parthkohli):

Yep. Now that's correct. And there's no need to 'answer' ;)

OpenStudy (anonymous):

@ParthKohli : i was talking about the length not area ....check it again

Parth (parthkohli):

Ah!